Suppose a monochromatic X-ray beam of wavelength 100 pm is sent through a Young’s double slit and the interference pattern is observed on a photographic plate placed 40 cm away from the slit. What should be the separation between the slits so that the successive maxima on the screen are separated by a distance of 0.1 mm?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
λ = 10 pm = 100 * 10^–12 m D = 40 cm = 40 * 10^–2 m β = 0.1 mm = 0.1 * 10^–3 m β = λD/d ⇒ d λD/β = 100* 10^-12 *40 * 10^-2/10^-3 *0.1 = 4 * 10^-7 m.